Phil wants to compare two means.

His sample statistics were GqgUxgEuncMzTxLa8CG4sqqCx_dLhEbmnHBYXTul1 = 22.7, s12 = 5.4, n1 = 9 and GqgUxgEuncMzTxLa8CG4sqqCx_dLhEbmnHBYXTul2 = 20.5, s22 = 3.6, n2 = 9.

Assuming equal variances, the test statistic is

Please provide solution calculations and excel solution
Answer #1

giren X = 22.7 512=54, n=9 & 22=2015, 522=3.6, n₂ =9 Assuming Equal variances, The test statisticisi t=(2, -Ñz) - 0 Where sos=21.06 -. t 22.7-2015 21.066 1 + = 2.2 21.06 X = 2.2 ll - 21.068014714 = 2.2 9.92.7684 = 0.2216 t = 0.2216) for the Here deg### By using Excel Spreadsheet:

1 Formula 22.7 20.5 9 s=21.06 2 xbaris 3 xbar2= 4 n1= 5 n2= s12= 7 s22= =((B4-1)*(B6*36)+(B5-1)*(B7*B7))/(B4-1+85-1) Test sta